OpenStudy (anonymous):

What is the distance between the following points? (-2 √7,10) and (4 √7,8)

OpenStudy (anonymous):

subtract Y2 - Y1 and X2 - X1

OpenStudy (anonymous):

i dont know how to subtract square roots

OpenStudy (anonymous):

if you have sqrt with same base you can just subtract the numbers out of it: 4 - (-2) = 4+2 = 6(sqrt{7})

OpenStudy (anonymous):

6sqrt7 +2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-2 becasue 8 -10 = -2

OpenStudy (anonymous):

hmm the answers ive got are a) 256 units b) sq.rt 352 units c)-6sq.rt7 + 2 units d) 16 units

OpenStudy (anonymous):

c would also be an aswer

OpenStudy (anonymous):

thought so. thanks so much